# Secure Information Embedding and Extraction in Forensic 3D Fingerprinting

### v1

@@ Abstract
The prevalence of 3D printing poses a significant risk to public safety, as any individual with internet access and a commodity printer is able to produce untraceable firearms, keys, counterfeit products, etc. To aid government authorities in combating these new security threats, several approaches have been taken to tag 3D-prints with identifying information. Known as fingerprints, this information is written into the object using various bit embedding techniques; examples include varying the height of the molten thermoplastic layers, and depositing metallic powder with different magnetic properties. Yet, the practicality of theses techniques in real-world forensic settings is hindered by the adversarial nature of this problem. That is, the 3D-printing process is out of reach of any law enforcement agencies; it is the adversary who controls all aspects of printing and possesses the printed object. To combat these threats, law enforcement agencies can regulate the manufacturing of 3D printers, on which they may enforce a fingerprinting scheme, and collect adversarially tampered remains (e.g., fragments of a broken 3D-printed firearm) during forensic investigation.
@@ 1 Introduction
3D printing brings about a revolution in consumption and distribution of goods. However, the prevalence of this technology exposes the public to various security threats. Any individual with internet access is able to produce untraceable firearms and counterfeits of IP-protected products (Figure. 1[1, 2, 3, 4]. To aid authorities and law enforcement agencies in combating these threats, fingerprinting 3D printed objects is one potential solution. In a fingerprinting scheme, a 3D-printer embeds uniquely traceable data into the printed object (e.g., timestamp, geoposition, printer ID, etc.), that can later be extracted to trace the perpetrator, e.g., if the object was confiscated or left at a crime scene.
@@
Various techniques for embedding bits in 3D printed objects have been proposed in the literature. For instance, embedding tags on the surface of the object by altering layer height [5], inserting cavities inside the object [6], altering the surface of the object by varying printing speed [7], embedding RFID-tags [8, 9, 10] or QR codes [11, 12], inserting acoustic barcodes using surface bumps [13], etc. For a broad introduction to the topic see [14, 15].
An orthogonal line of work in the forensic 3D-printing literature proposes to identify printers from 3D-prints using machine-learning methods, based on manufacturing imperfections of printer components reflected on the print results [16]. However, whether the proposed method provides enough resolution for a large number of 3D printers remains unknown. Moreover, [16] focuses on identifying the 3D printer from the printed object, and does not provide a method for embedding arbitrary bits. These bits, which may represent geoposition or timestamp, are unknown at the time of recording the printer’s imperfections, and are crucial for forensic applications.
The practicality of any such bit embedding technique, however, is hindered by a fundamental obstacle, i.e., the adversary controls the printing. Specifically, the adversary has full control over the entire printing process, and may manipulate the prints after committing a crime. For example, the adversary may bypass the information embedding software installed on the printer control board and print a fingerprint-free firearm. The firearm may also be broken after committing the crime, making the embedded information unrecoverable.
In contrast, law enforcement agencies are restricted to merely regulate the manufacturing and distribution of commercial 3D printers. Once the printer is handed over to the (potential) adversary, the agencies can hardly monitor its usage, and can only attempt to recover information from adversarially tampered 3D-prints during forensic crime investigation. Therefore, a practical information embedding scheme must guarantee successful embedding and successful extraction of a reasonable number of bits, even in circumstances where printing is out of reach of law enforcement authorities, and even in circumstances where the 3D-prints are adversarially damaged (Figure 2).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionpt3D PrinterptLaw Enforcement AgencypttamperptprintptmanufactureptCompaniesptAdversary
@@
Indeed, any embedded bits can be stripped-off from the 3D model file before printing, and any tag can be stripped-off using household items, as the printed object itself could be melted, deformed, or simply broken apart. As such, the aforementioned methods fail in the face of a knowledgeable adversary that tampers with the printing process and the prints, as they rely on the integrity of both the information embedding process and the geometric shape of the printed object.
To this end, we propose SIDE, a secure information embedding and extraction framework that is robust against a knowledgeable adversary that controls the printing process and may tamper with the printed object. Specifically, our contribution in this work is two-fold:
  • Secure information embedding: SIDE incorporates a Trusted Execution Environment (TEE) into the 3D printing process in order to protect its integrity. Consequently, the information embedding procedure is guaranteed to succeed even if the adversary is able to compromises the printer.
  • Secure information extraction: SIDE employs coding-theoretic techniques for correct information extraction from the 3D-prints even if their geometric shapes are adversarially damaged. As a result, the law enforcement agency is guaranteed to extract information correctly.
@@ 2 Background and Related Work
3D Printing, commonly referred to as Additive Manufacturing (AM), has emerged as a revolutionary technology with profound implications across various industries. In contrast to the traditional subtractive manufacturing during which materials are consecutively removed from the workpiece, 3D printing refers to an additive process of creating a physical object and is typically done by laying down many thin layers of material (such as liquid plastic or powdered metal) in succession. Indeed, the manufacturing process starts with “slicing” the to-be-manufactured 3D model into discrete horizontal layers/slices using a software called slicer. Each layer represents a planar cross-section of the model along the printing direction (the $z$-direction).
Numerous technologies have been developed for 3D printing. By and large, they differ by which material is in use and how layers are formed. A layer can either be formed by using a nozzle that deposits molten thermoplastics while shifting back and forth on a surface, e.g. Fused Deposition Modeling (FDM), by depositing a layer of liquid polymers and curing it by light exposure, e.g., Stereolithography (SLA), or by depositing powdered material over the surface and binding it using high-energy laser beams, e.g., Selective Laser Sintering (SLS).
In this research, we focus on the FDM technology for its prevalence in commodity 3D printers (compared with SLS), as well as the profound mechanical strength offered to the printed objects (compared with SLA). In an FDM-based printer, the slicer operates by converting the layers to a G-code, which is a series of commands for the printer’s hardware. Such commands may include the nozzle movement along the $x$$y$, and $z$ axes, the extrusion of material, the nozzle/bed temperature, and other specific aspects based on the printer’s capabilities.
It is worth noting that a G-code command specifies only the expected action of the printer hardware in a relatively high level, while the low-level implementation is not addressed. For example, the command
@@
merely instructs the printer to move its nozzle to $(98.302,96.831)$ in the $x$-$y$ coordinate and simultaneously extrude $15.796$ millimeter of molten thermoplastics. Completing such an operation requires a series of discrete angular-step rotations of multiple stepper motors, which are directly operated by the printer firmware.
Indeed, the firmware plays a crucial role in bridging the G-code and printer hardware. In spite of commands for heaters and other applicable components, the firmware translates the G-code to signals that control the kinematics of the printer’s actuators, e.g., stepper motors. This translation task is non-trivial, and has a significant impact on print quality. For instance, a sudden jerk of the printer nozzle may lead to uneven deposition of print material, which jeopardizes the final print quality or even fails the print. In contrast, a nozzle movement with smooth velocity change is generally preferred. Therefore, the print quality may be significantly improved with an updated firmware, and the same G-code may be executed on printers with completely different builds.
@@ 2.2 Existing 3D Fingerprinting Methods
Similar ideas have been implemented successfully in several recent works. Ref. [5] varies the thickness of each layer across several adjacent layers to create a matrix of bits that is visible to the naked eye on the surface of the object. Parity bits were then added to resolve reliability issues in some cases, and additional noise patterns were discussed, such as orientation issues and sanding. In the method LayerCode [17], variations in color and thickness were used to embed a barcode on the surface of the printed object, that can be retrieved using a smartphone camera.
@@
An orientation-based method has been implemented in [18], where the authors print a reference layer that is circularly grooved by a sequence with low auto-correlation. Data is embedded in all other layers by the respective angle of the layer to the reference layer; this enables encoding with alphabet size larger than two.
@@
Various other creative ideas have been explored. Examples include embedding information-carrying cavities within the object [19, 20], water-marking the 3D-mesh of the surface of the object [21], inserting RFID tags [8], inserting a series of notches which create an acoustic barcode when tapped [13], etc. In the data extraction phase, most existing methods rely on an RGB camera, a 3D scanner, or an ordinary scanner. Future technologies however, such as the ones in Figure 3, might require an industrial CT scanner.
@@
However, we argue that none of these approaches is suitable for forensic applications. First, they all implicitly assume that a mechanism for correct information embedding is in place. Such a mechanism is crucial since in most scenarios the adversary owns the printer and/or the file, and might potentially remove the embedded bits altogether. Second, none of the methods is provably resilient to adversarial tampering; they can be easily breached by a knowledgeable adversary that can scrape the object or break it apart.
@@
A different line of works applies machine-learning based methods to identify the intrinsic fingerprint which results from printer hardware imperfections (PrinTracker [16]). The variations of printer components arising from manufacturing errors are essentially inevitable, and create perceivable differences between objects made from different 3D-printers, even if they are printed from the same 3D model. These differences form a unique printer fingerprint, and can be extracted from the printed objects with commodity scanners.
@@
Unlike the aforementioned techniques, PrinTracker is a fingerprinting solution which considers the adversarial nature of forensic applications, as well as potential attacks. This is different from the information embedding solution proposed in our work. Although information embedding can be used for fingerprinting purpose, it is not limited to it. The proposed solution allows for arbitrary bit embedding, which enables a variety of forensic applications. For example, certain applications may require the embedding of timestamp or geoposition of the printing into the printed object in order to check whether the printer has been misused in unauthorized time and location. Further, it remains to be shown if the extraction, storage, and identification of such imperfections scales well to a large number of printers.
@@ 2.3 Coding methods
Similar coding problems of information extraction from fragments have been previously studied in the literature, motivated by applications in DNA storage. In particular, several variations of the so-called torn-paper channel were studied in [22, 23, 24, 25, 26], where [22, 23, 24] focused on a probabilistic error model which is incompatible with our adversarial setting, and [25] studied an adversarial error model in which fragment are restricted in length. Most closely related, [26] studied $t$-break codes, which coincide with our problem definition (given shortly). They provided a theoretical analysis of the fundamental limits, and an (almost) matching code construction. However, the scheme described in [26] involves random encoding and is only effective for a very large number of embedded bits. Our work herein addresses the same model of [26], yet at a much simplified framework which can be easily employed in practical settings, albeit below the maximum theoretical information rate.
@@ 3 Threat Model
We consider an knowledgeable adversary who wishes to commit a crime. The adversary is unable to obtain a certain required instrument, say, fails to pass background checks when attempting to purchase a firearm, or has insufficient financial means. Hence, the adversary manufactures the instruments independently, by downloading a 3D model file and feeding it to a commercial grade 3D-printer bought from a local electronics retailer.
We assume that the 3D printer, which does not require a background check to purchase, holds a binary string $\mathbf{w}$. The content of such string may include a unique printer ID, the time and geoposition of the printing process, the identity of the file, etc, which aid the authorities to trace the perpetrator111 In the upcoming experiments we focus solely on printer ID. The secure procurement and embedding of additional real-time identifying information is left for future studies. .
The string $\mathbf{w}$ is embedded in every object printed by that printer, in an arbitrary fashion [^2] Our methods are not inherently restricted to any particular bit embedding technique, as long as those bits are manifested as physical elements on the printing direction along the object. For the sake of experiments, we embed bits by varying layer heights.. We assume that the actuators (e.g., stepper motors) and other electronics (e.g., heater) of the 3D-printer are connected to a hardware platform (e.g., a control board) which supports a trusted execution environment (TEE). The TEE provides strong isolation on processor, memory, and peripherals (necessary for connecting the components). The connection from the board to the electronic components is also free of malicious tampering.
Meanwhile, the adversary is able to compromise the rich execution environment (REE) on the control board, including all necessary software components and the hosting operation system. Then, the adversary prints the object, commits the crime, and mechanically damages the printed tool, breaking it into at most $t$ fragments along the printing direction (i.e., the “$z$” direction). The reasoning behind this assumption stems from the following observations.
@@
    3D printers operate in a layer-by-layer fashion, and the 3D-prints have the lowest mechanical strength along the printing direction. If the break crosses multiple layers, then the irregular geometry of the broken section is likely to enable correct reassembly.
@@
    The parameter $t$ reflects time constraints, access to instruments, or simply physical strength limitations, which prevent the adversary from breaking the object to arbitrarily many fragments. Hence, the parameter $t$ is expected to be small with respect to the number of embedded bits.
We assume the fragments of the broken instrument are confiscated by a law enforcement agency after the crime has been committed. The agency is able to read the embedded bits from the fragments by examining the height of each layer. Examples of equipment for this task includes microscopes and computed tomography (CT) devices. However, since the bit strings are extracted from at most $(t+1)$ fragments, the agency is not aware of their correct order.
4 SIDE Overview
SIDE incorporates coding-theoretic techniques with hardware isolation mechanism for secure information embedding and secure information extraction.

### v3

@@ Abstract
The prevalence of 3D printing poses a significant risk to public safety, as any individual with internet access and a commodity printer is able to produce untraceable firearms, keys, counterfeit products, etc. To aid government authorities in combating these new security threats, several approaches have been taken to tag 3D-prints with identifying information. Known as a fingerprint, this information is written into the object using various bit embedding techniques; examples include varying the thickness of the molten thermoplastic layers, and depositing metallic powder with different magnetic properties. Yet, the practicality of these techniques in real-world forensic settings is hindered by the adversarial nature of this problem. That is, the 3D-printing process is out of reach of any law enforcement agencies; it is the adversary who controls all aspects of printing and possesses the printed object. To combat these threats, law enforcement agencies can regulate the manufacturing of 3D printers, on which they may enforce a fingerprinting scheme, and collect adversarially tampered remains (e.g., fragments of a broken 3D-printed firearm) during forensic investigation.
@@ 1 Introduction
3D printing brings about a revolution in consumption and distribution of goods. However, the prevalence of this technology exposes the public to various security threats. Any individual with a commercial 3D printer and internet access is able to produce untraceable firearms and counterfeits of IP-protected products (Figure 1[CNET, Yahoo, McNulty, ENGADGET]. To aid authorities and law enforcement agencies in combating these threats, fingerprinting 3D printed objects is one potential solution. In a fingerprinting scheme, a 3D-printer embeds uniquely traceable data into the printed object (e.g., timestamp, geoposition, printer ID, etc.), that can later be extracted to trace the perpetrator, e.g., if the object was confiscated or left at a crime scene.
@@
Various techniques for embedding bits in 3D printed objects have been proposed in the literature. For instance, embedding tags on the surface of the object by varying layer thickness [LayerThickness], altering layer material [salas2022embedding], inserting cavities inside the object [Cavities], altering the surface of the object by varying printing speed [elsayed2021information], embedding RFID-tags [voris2017three, RFID2, RFID1] or QR codes [wei2018embedding, chen2019embedding], inserting acoustic barcodes using surface bumps [Acoustic], etc. For a broad introduction to the topic see [usama2022embedding, elsayed2023information]. An orthogonal line of work in the forensic 3D fingerprinting literature, which does not involve embedding nor extraction of bits, is based on identifying printers from their 3D-prints. This is done using machine-learning methods, based on manufacturing imperfections of printer components reflected on the print results [Imperfections].
The practicality of any such bit embedding technique, however, is hindered by a fundamental obstacle, i.e., the adversary controls the printing. First, the adversary has full control over the entire printing process. This gives the adversary the ability to manipulate information embedding, e.g., the adversary may bypass the information embedding module installed on the printer and print a fingerprint-free firearm. Second, the adversary is in possession of the 3D prints, which may be broken after committing the crime, making the embedded information unrecoverable.
@@
On the other hand, law enforcement agencies are restricted to merely regulate the manufacturing and distribution of commercial 3D printers. Once the printer is handed over to the (potential) adversary, the agencies can hardly monitor its usage, and can only attempt to recover information from adversarially tampered 3D-prints during forensic crime investigation (Figure 2).
Indeed, any embedded bits can be stripped-off from the 3D model file before printing, and any tag can be stripped-off using household items, as the printed object itself could be melted, deformed, or simply broken apart. As such, the aforementioned methods fail in the face of an adaptive adversary [^1] An adaptive adversary in a knowledgeable attacker that is fully aware of the information embedding/extracting scheme. that tampers with the printing process and the prints, since they rely on the integrity of both the information embedding process and the geometric shape of the printed object. Therefore, a practical fingerprinting scheme must overcome the following two key challenges under the circumstances in which both the printing and the prints are out of reach of law enforcement authorities.
Challenge 1 (Embedding): Guarantee successful information embedding, even in the face of an adaptive adversary that is fully aware of, and is able to interfere with, the embedding process.
Challenge 2 (Extraction): Guarantee successful information extraction, even in circumstances where the 3D-prints are adversarially tampered.
To this end, we design and implement SIDE, a Secure Information EmbeDding and Extraction framework that is robust against an adaptive adversary that controls the printing process and may tamper with the printed object. Specifically, our contribution in this work is three-fold:
  • Secure information embedding: We incorporate SIDE with a Trusted Execution Environment (TEE) into the 3D printing process in order to protect its control-flow integrity. Consequently, the information embedding procedure is guaranteed to succeed even if the adversary is able to compromises the printer’s OS.
  • Secure information extraction: We develop break-resilient codes (BRC), which are coding-theoretic techniques for information extraction from the 3D-prints, and incorporate them with SIDE. We develop corresponding decoding algorithms, and formally prove that they correctly extract the information even if the geometric shapes are adversarially broken to at most a certain number of parts; the maximum supported number of parts serves as a security parameter. As a result, the law enforcement agency is guaranteed to extract the information correctly.
  • Implementation: We implemented and tested a prototype of SIDE. On the secure information embedding side, we realize SIDE on a Creality Ender 3 3D printer controlled by a Raspberry Pi 3B board, with OP-TEE V3.4 support enabled in its ARMv8-A architecture to protect the control-flow integrity of the printing process. On the secure information extraction side, we implemented efficient BRC encoder and decoder. The encoded information is embedded in the printed object by varying layer thickness, and is read by a Leica S9D microscope. In multiple experiments, we printed objects with embedded information, broke them apart, and managed to read out the embedded information successfully.
@@ 2 Background and Related Work
3D Printing, commonly referred to as Additive Manufacturing (AM), has emerged as a revolutionary technology with profound implications across various industries. In contrast to the traditional subtractive manufacturing during which materials are consecutively removed from the workpiece, 3D printing refers to an additive process of creating a physical object and is typically done by laying down many thin layers of material in succession.
Numerous technologies have been developed for 3D printing. By and large, they differ by which material is in use and how layers are formed. A layer can either be formed by using a nozzle that deposits molten thermoplastics while shifting back and forth on a surface, e.g. Fused Deposition Modeling (FDM), by depositing a layer of liquid polymers and curing it by ultraviolet light exposure, e.g., Stereo lithography Appearance (SLA), or by depositing powdered material over the surface and binding it using high-energy laser beams, e.g., Selective Laser Sintering (SLS).
The manufacturing process starts with generating G-code commands from the user-input 3D model file. In detail, the slicer software operates by converting the model into discrete 2D diagrams, each represents a planar cross-section of the model along the printing direction at a certain height. Then, the slicer creates a series of G-code commands to instruct the printer about how to produce the corresponding layers of the diagram sequence.
In this research, we focus on the FDM technology for its prevalence in commodity 3D printers (compared with SLS), as well as the profound mechanical strength offered to the printed objects (compared with SLA). As a result, the G-code commands for FDM printers may include the nozzle movement along the $x$$y$, and $z$ axes, the extrusion of material, nozzle/bed temperature, and other specific aspects based on the printer’s capabilities.
Figure 3: Illustration of the printing process. At first, the user feeds the 3D model model.obj to the slicer software and obtains model.gcode, which involves a series of commands for the printer. The printer firmware then translates these high-level commands to low-level stepping events and other hardware control instructions (stpr_evnt), which are then fed to the printer hardware for the physical printing process. To guarantee secure information embedding, a fingerprinting scheme must protect the control-flow integrity of this process.
A typical FDM printer involves four stepper motors, which are actuators that rotate in discrete angular steps of constant degree. Three of the motors control the nozzle movement in the Cartesian space, and one is responsible for filament extrusion. However, a G-code command specifies only the expected action of the printer hardware in a relatively high level, while the low-level implementation is not addressed. For example, the command
@@
merely instructs the printer to move its nozzle from its current position to location $(98.302,96.831)$ in the $x$-$y$ coordinate (with its position relative to the $z$ axis unchanged), and simultaneously extrude $15.796$ millimeter of molten thermoplastics. Completing this operation requires a series of stepping events, and each of them defines the exact timing and direction to trigger one of the four stepper motors for a single angular step.
To this end, the printer firmware plays a crucial role in bridging the G-code and printer hardware. In addition to commanding heaters and other applicable components, the firmware translates the G-code commands to a series of stepping events for the actuators, and the execution of them concludes the printing process. See Figure 3 for a complete illustration of the printing process.
It is worth noting that the translation task is non-trivial, and has a significant impact on print quality. For instance, a sudden jerk of the printer nozzle may lead to uneven deposition of print material, which jeopardizes the final print quality or even fails the print. In contrast, a nozzle movement with smooth velocity change is generally preferred. Therefore, the print quality may be significantly improved with an updated firmware, and the same G-code may be executed on printers with completely different builds.
@@ 2.2 Existing 3D Fingerprinting Methods
Similar ideas have been implemented successfully in several recent works. Delmotte et al. vary the thickness of each layer across several adjacent layers to create a matrix of bits that is visible to the naked eye on the surface of the object [LayerThickness]. Parity bits were then added to resolve reliability issues in some cases, and additional noise patterns were discussed, such as orientation issues and sanding. In the method LayerCode [LayerCode], variations in color and thickness were used to embed a barcode on the surface of the printed object, that can be retrieved using a smartphone camera.
@@
An orientation-based method has been implemented in [hou20153d], where the authors print a reference layer that is circularly grooved by a sequence with low auto-correlation. Data is embedded in all other layers by the respective angle of the layer to the reference layer; this enables encoding with alphabet size larger than two.
@@
Various other creative ideas have been explored, including embedding information-carrying cavities within the object [willis2013infrastructs, li2017aircode], water-marking the 3D-mesh of the surface of the object [cho2006oblivious], inserting RFID tags [voris2017three], inserting a series of notches which create an acoustic barcode when tapped [Acoustic], etc. In the data extraction phase, most existing methods rely on an RGB camera, a 3D scanner, or an ordinary scanner. Future technologies however, such as the ones in Figure 4, might require an industrial CT scanner.
@@
However, we argue that none of these approaches is suitable for forensic applications. First, they all implicitly assume that a mechanism for correct information embedding is in place. Such a mechanism is crucial since in most scenarios the adversary owns the printer and/or the file, and might potentially remove the embedded bits altogether. Second, none of the methods is provably resilient to adversarial tampering; they can be easily breached by an adaptive adversary that can scrape the object or break it apart.
@@
PrinTracker [Imperfections] represents a different line of works, which apply machine-learning based methods to identify the intrinsic fingerprint which results from printer hardware imperfections. The variations of printer components arising from manufacturing errors are essentially inevitable, and create perceivable differences between objects made from different 3D-printers, even if they are printed from the same 3D model. These differences form a unique printer fingerprint, and can be extracted from the printed objects with commodity scanners.
@@
Unlike the aforementioned techniques, PrinTracker does not embed bits into the object; although it also considers the adversarial nature of forensic applications, as well as potential attacks, it is different from the information embedding solution proposed in our work. Further, it remains to be shown if the extraction, storage, and identification of such imperfections scales well to a large number of printers. Our proposed solution allows for embedding arbitrary bits, which enables a variety of forensic applications, including but not limited to fingerprinting. For example, certain applications may require the embedding of timestamp or geoposition of the printing into the printed object in order to check whether the printer has been misused in unauthorized time and location.
2.3 Trusted Execution Environment
Trusted Execution Environments (TEEs) have emerged as a solution to the problem of protecting sensitive and critical operations in modern computing systems. With hardware and/or software isolation mechanisms, a TEE provides a secure execution space for the processes running within, which are commonly referred as Trusted Applications (TAs). The TAs, as well as their assets, cannot be accessed or tampered by unauthorized codes from the outside world, i.e., the main operating system which is commonly referred to as the Rich Execution Environment (REE) or normal world.
Hardware implementations of TEE include ARM TrustZone, Intel Software Guard Extensions (SGX), and RISC-V Keystone. Meanwhile, software development kits (SDKs) for TEE are also available [optee, sgxsdk] for the ease of TA developments. Together, TEE have hosted applications across numerous domains, including access control [krawiecka2018safekeeper, schwarz2020seng, djoko2019nexus], could computing [baumann2015shielding, schuster2015vc3, poddar2018safebricks, hunt2020telekine, mo2021ppfl], and real-time systems [alder2021aion, wang2022rt, pinto2019virtualization].
@@ 2.4 Coding Methods
Secure information extraction is essentially a problem of communicating in the presence of (potentially adversarial) noise. Such problems are studied in the information-theoretic literature under the title of coding theory [roth2006coding]. A typical problem setup in coding theory includes a sender (e.g., the printer), which would like to send a message (e.g., printer ID) to a receiver (e.g., law-enforcement). This must be accomplished successfully even if an adversary (e.g., a criminal) injects adversarial noise that is limited by some security parameter.
Coding problems of information extraction from fragments have been previously studied in the literature, motivated by applications in DNA storage. In particular, several variations of the so-called torn-paper channel were studied in [shomorony2020communicating, shomorony2021torn, ravi2021capacity, bar2023adversarial, wang2023breakresilient], where [shomorony2020communicating, shomorony2021torn, ravi2021capacity] focused on a probabilistic error model which is incompatible with our adversarial setting, and [bar2023adversarial] studied an adversarial error model in which fragment are restricted in length. Most closely related, [wang2023breakresilient] studied $t$-break codes, which coincide with our problem definition (given shortly). They provided a theoretical analysis of the fundamental limits, and an (almost) matching code construction. However, the scheme described in [wang2023breakresilient] involves random encoding and is only effective for a very large number of embedded bits. Our work herein addresses the same model of [wang2023breakresilient], yet at a much simplified framework which can be easily employed in practical settings, albeit below the maximum theoretical information rate.
@@ 3 Threat Model
We consider an adaptive adversary, i.e., one who is fully aware of the proposed embedding/extracting scheme. The adversary wishes to commit a crime, but is unable to obtain a certain required instrument, say, fails to pass background checks when attempting to purchase a firearm, or has insufficient financial means. Hence, the adversary manufactures the instruments independently, by downloading a 3D model file from the internet and feeding it to a commercial grade 3D-printer bought from a local electronics retailer.
We assume that the 3D printer does not require a background check to purchase. At the time of printing, the printer composes a binary string $\mathbf{w}$ to be embedded within the object. The content of such string may include a unique printer ID, the time and geoposition of the printing process, the identity of the file, etc, which aid the authorities to trace the perpetrator222 In the upcoming implementations and experiments we focus solely on printer ID. The secure procurement and embedding of additional real-time identifying information is left for future studies. .
The string $\mathbf{w}$ is embedded in every object printed by that printer using physical elements on the printing direction, e.g., altering layer thickness or material. We assume that the actuators (e.g., stepper motors) and other electronics (e.g., heater) of the 3D-printer are connected to a hardware platform (e.g., a control board) which supports a trusted execution environment (TEE). The TEE provides strong isolation on processor, memory, and peripherals. Specifically, the General-Purpose Input/Output (GPIO) port connecting to the printer hardware is driven by a secure program that resides inside the TEE, and the connection from the board to the hardware is also free of malicious tampering.
Meanwhile, the adversary is able to compromise the rich execution environment (REE) on the control board, including all necessary software components and the hosting operation system. Then, the adversary prints the object, commits the crime, and mechanically damages the printed tool, breaking it into pieces.
Figure 5: Possible types of break that cross multiple layers. In either case, the assembly of pieces may be inferred from their geometries.
Figure 6: Fragments of a transmission shaft that is broken along the $z$ direction. The correct assembly (i.e., order) cannot be inferred from their geometries.
Breaks fall into one of two categories, depending on whether it crosses multiple layers or not. We intentionally disregard the former category of breaks, as it is likely to cause an irregular geometry of the broken section, which enables correct reassembly of the pieces. Moreover, the identical layer-thickness pattern of two matching fragments provide enough hints to couple them together (See Figure 5 for illustrative examples).
Instead, we are particularly interested in the latter type of breaks, which occur along the printing direction and separate adjacent layers of the printed object. Our preference stems from the following observations.
@@
    3D printers operate in a layer-by-layer fashion, and as a result, the printed objects have the lowest mechanical strength along the printing direction, making this type of breaks more frequent.
@@
    This type of breaks reveals limited information on the geometry of the tool that allows for reassembly (See Figure 6 for an illustrative example).
We assume that the pieces of the broken tool are confiscated by a law enforcement agency after the crime has been committed. The agency is able to reassemble breaks of the first category (Figure 5), but not of the second (Figure 6). This results in at most $t+1$ fragments caused by at most $t$ breaks of the second category, whose correct reassembly is unknown, where $t$ is a security parameter that reflects time constraints, access to instruments, or physical strength limitations. The parameter $t$ is crucial for the analysis of the system, since it prevents the adversary from breaking the object into arbitrarily many fragments, thus eliminate all chances of extracting the fingerprint.
Having obtained these $t+1$ unordered fragments, the agency reads the embedded bits from the fragments by examining the bit-embedding physical elements on the printing direction. The reading mechanism requires the use of additional technology such as a smartphone camera, a microscope, or a computed tomography (CT) device, depending on the thickness of the layers, the complexity of the object, the outside finish of the object (e.g., coating), and the external damage done by the adversary (e.g., scraping).
Finally, we emphasize that we do not rely on correct reassembly of the $t+1$ unordered fragments of the object, and base our secure information extraction techniques solely on the information content of the fragments. That is, the decoder (i.e., law-enforcement) is given a collection of $t+1$ unordered fragments, which cannot be assembled based on their geometric shape, and needs to recover the information successfully. Recovery must be successful in all cases in which the adversary breaks the object apart at most $t$ times, in arbitrary positions along the $z$ direction.
4 Secure Information Embedding
4.1 Preliminary and overview
To provide secure embedding, one must protect the control-flow integrity of the printing process shown in Figure 3. That is, the printer hardware must receive the instructions generated by the firmware. Meanwhile, the firmware must receive the G-code generated by the slicer, with the desired information embedded within.
However, the integrity of this control flow can be easily jeopardized by an adversary with full control over the REE of the printer control board. For example, this adversary may bypass the slicing process by feeding to the firmware a G-code file that was generated independently (i.e., not by the designated slicer), and does not contain the embedded information. Alternatively, the adversary may control the printer hardware using a modified firmware, in order to similarly alter the information embedding process.
To address the aforementioned issues, SIDE employs a Trusted Execution Environment (TEE) to protect the control-flow integrity of the printing process; such a design is motivated by the resemblance between 3D printing and the compilation process of a program (e.g., a C program). Indeed, similar to human-readable C code which must be converted to assembly/binary, the 3D models must be converted to G-code which the printer firmware can recognize.